# Monte Carlo sampling with integrator snippets

### v1

@@ 2.2 Justification outline
First for $(n,k)\in\llbracket 0,P\rrbracket\times\llbracket 0,T\rrbracket$ let $\psi_{n,k}\colon\mathsf{Z}\rightarrow\mathsf{Z}$ be measurable and invertible mappings, define $\mu_{n,k}({\rm d}z):=\mu_{n}^{\psi_{n,k}^{-1}}({\rm d}z)$, i.e. the distribution of $\psi_{n,k}^{-1}(z)$ when $z\sim\mu_{n}$. It is worth pointing out that invertibility of these mappings is not necessary, but facilitates interpretation throughout, as illustrated by the last statement about the distribution of $\psi_{n,k}^{-1}(z)$. Earlier we have focused on the scenario where $\psi_{n,k}=\psi_{n}^{k}$ for an integrator $\psi_{n}\colon\mathsf{Z}\rightarrow\mathsf{Z}$, but this turns out not to be a requirement, although it is our main motivation. Useful applications of this general perspective can be found in Subsection 2.3. Introduce the probability distributions on $\big{(}\mathbb{\llbracket}0,T\rrbracket\times\mathsf{Z},\mathscr{P}(\mathbb{\llbracket}0,T\rrbracket)\otimes\mathscr{Z}\big{)}$
@@
which we may refer to as a mixture, for $n\in\llbracket 0,P\rrbracket$. Note that the conditional distribution is
@@
For $n\in\llbracket 0,P\rrbracket$ assume the existence of a $\sigma$-finite dominating measure $\upsilon\gg\mu_{n}$, therefore implying the existence of a density $\mu_{n}(z):={\rm d}\mu_{n}/{\rm d}\upsilon(z)$; $\upsilon$ could be the Lebesgue measure. Assuming that $\upsilon$ is $\psi_{n,k}$-invariant, or “volume preserving”, then Lemma 32 implies, for $k\in\llbracket 0,T\rrbracket$,
@@
When $\upsilon$ is the Lebesgue measure and $\{\psi_{n,k},k\in\llbracket P\rrbracket\}$ are not volume preserving, additional multiplicative Jacobian determinant-like terms may be required (see Lemma 32 and the additional requirement that $\{\psi_{n,k},k\in\llbracket P\rrbracket\}$ be differentiable). However this extra term may be more complex and require application specific treatment; adoption of the measure theoretic notation circumvents such peripheral considerations at this stage, which can be ignored until actual implementation of the algorithm.
A central point throughout this paper is how samples from $\mu_{n}$ can be used to obtain samples from the marginal $\bar{\mu}_{n}$ and vice-versa, thanks to the mixture structure relating the two distributions. Naturally, given $z\sim\mu_{n}$, sampling $k\sim\mathcal{U}\big{(}\llbracket 0,T\rrbracket\big{)}$ and returning $\psi_{n,k}^{-1}(z)$ yields a sample from the marginal $\bar{\mu}_{n}$. Now assuming $z\sim\bar{\mu}_{n}$ and then sampling $k\sim\bar{\mu}_{n}(\cdot\mid z)$ naturally yields $(k,z)\sim\bar{\mu}_{n}$ and hence intuitively $\psi_{n,k}(z)\sim\mu_{n}$. We now formally establish a more general result concerned with the estimation of expectations with respect to $\mu_{n}$ from samples from $\bar{\mu}_{n}$. For $f\colon\mathsf{Z}\rightarrow\mathbb{R}$ $\mu_{n}$-integrable and $k\in\llbracket 0,T\rrbracket$, a change of variable (see Theorem 31) yields
@@
which formalizes the fact, at first sight of limited interest, that $\mu_{n}^{\psi_{n,k}^{-1}}$ is the distribution of $\psi_{n,k}^{-1}(z)$ for $z\sim\mu_{n}$ and therefore $\psi_{n,k}\circ\psi_{n,k}^{-1}(z)\sim\mu_{n}$. The relevance of this remark comes from the identity
@@
$=\sum_{k=0}^{T}\int f\circ\psi_{n,k}(z)\bar{\mu}_{n}(k,{\rm d}z)$ (10)
=$=\int\Big{\{}\sum_{k=0}^{T}f\circ\psi_{n,k}(z)\bar{\mu}_{n}(k\mid z)\Big{\}}\bar{\mu}_{n}({\rm d}z),$ (11)
which implies that samples from the mixture $\bar{\mu}_{n}$ can be used to unbiasedly estimate $\mu_{n}(f)$ thanks to an appropriate weighted average along the snippet $k\mapsto f\circ\psi_{n,k}(z)$. Using $f=\mathbf{1}_{A}$ for $A\in\mathscr{Z}$ formally establishes the earlier claim that if $(k,z)\sim\bar{\mu}_{n}$ then $\psi_{n,k}(z)\sim\mu_{n}$. Note that, as suggested by Example 2, construction of the estimator will only require evaluations of the density $\mu_{n}$ and function $f$ at $z,\psi_{n,1}(z),\psi_{n,2}(z),\ldots,\psi_{n,T}(z)$.
We now turn to the description of an SMC algorithm targeting $\big{\{}\bar{\mu}_{n},n\in\llbracket 0,P\llbracket\big{\}}$, Alg. 3, and then establish that it is probabilistically equivalent to Alg. 2 in a sense made precise below. With $z=(x,v)$ for $n\in\llbracket P\rrbracket$ we introduce the following mutation kernel
@@
where we note that the refreshment kernel has the property that $\mu_{n-1}R_{n}=\mu_{n-1}$. One can show that the near optimal kernel $\bar{L}_{n-1}\colon\mathsf{Z}\times\mathscr{Z}\rightarrow[0,1]$ is given for any $(z,A)\in\mathsf{Z}\times\mathscr{Z}$ by
and is well defined $\bar{\mu}_{n-1}\bar{M}_{n}$-a.s. (see Lemma 4, given at the end of this subsection) and satisfies the property that for any $A,B\in\mathscr{Z}$
that is $\bar{L}_{n-1}(z,A)$ is a conditional probability of $A$ given $z$ for the joint probability distribution $\bar{\mu}_{n-1}\otimes\bar{M}_{n}$. With the assumption $\mu_{n-1}\gg\bar{\mu}_{n}$, from Lemma 4, the SMC sampler importance weights at step $n\in\llbracket P\rrbracket$ are
@@
that is for $f\colon\mathsf{Z}\rightarrow\mathbb{R}$ such that $\bar{\mu}(|f|)<\infty$,
$\int f(z^{\prime})\frac{{\rm d}\bar{\mu}_{n}}{{\rm d}\mu_{n-1}}(z^{\prime})\bar{\mu}_{n-1}\otimes\bar{M}_{n}\big{(}{\rm d}(z,z^{\prime})\big{)}=\bar{\mu}(f)\,.$
@@
Notice that we assume here $\psi_{0}={\rm Id}$, and hence that $\bar{\mu}_{0}=\mu_{0}$, and that the weights $w_{n,k}$ appear as being computed twice in step 3 and step 14 when evaluating the resampling weights at the previous iteration, for the only reason that it facilitates exposition. The identities (11) and (9) suggest, for $n\in\llbracket P\rrbracket$ the estimator of $\mu_{n}(f)$, for $f\colon\mathsf{Z}\rightarrow\mathbb{R}$ $\mu_{n}$-integrable,
@@
a form of Rao-Blackwellization implying lower variance for $\hat{\mu}_{n}(f)$ while the two estimators share the same bias. The result is in fact stronger since it states that the estimators are convex ordered, a property which we however do not exploit further here. Interestingly a result we establish later in the paper, Proposition 15, suggests that the variance $\check{\mu}_{n}(f)$ is smaller than that of the standard Monte Carlo estimator assuming $N$ samples $z_{n}^{(i)}\overset{{\rm iid}}{\sim}\mu_{n}$, due to the control variate nature of integrator snippets estimators.
An interesting point is that computation of the weight $\mu_{n}\circ\psi_{n,k}/\bar{\mu}_{n}$ only requires knowledge of $\mu_{n}$ up to a normalizing constant, that is the estimator is unbiased if $\check{z}_{n}^{(i)}\sim\bar{\mu}_{n}$ for $i\in\llbracket N\rrbracket$ even if $\mu_{n}$ is not completely known, while the estimator (7) will most often require self-normalisation, hence inducing a bias.
We now provide the probabilistic argument justifying Alg. 2 and the shared notation $\{z_{n}^{(i)},i\in\llbracket N\rrbracket\}$ in Alg. 2 and Alg. 3.

### v2

@@ 2.2 Justification
First for $(n,k)\in\llbracket 0,P\rrbracket\times\llbracket 0,T\rrbracket$ let $\psi_{n,k}\colon\mathsf{Z}\rightarrow\mathsf{Z}$ be measurable and invertible mappings, define $\mu_{n,k}({\rm d}z):=\mu_{n}^{\psi_{n,k}^{-1}}({\rm d}z)$, i.e. the distribution of $\psi_{n,k}^{-1}(z)$ when $z\sim\mu_{n}$. It is worth pointing out that invertibility of these mappings is not required theoretically, but important practially and facilitates interpretation throughout. Earlier we have focused on the scenario where $\psi_{n,k}=\psi_{n}^{k}$ for an integrator $\psi_{n}\colon\mathsf{Z}\rightarrow\mathsf{Z}$, but this turns out not to be a requirement, although it is our main motivation. Useful applications of this general perspective can be found in Subsection 2.4. Introduce the probability distributions on $\big{(}\mathbb{\llbracket}0,T\rrbracket\times\mathsf{Z},\mathscr{P}(\mathbb{\llbracket}0,T\rrbracket)\otimes\mathscr{Z}\big{)}$
@@
which we may refer to as a mixture. Note that the conditional distribution is
@@
For $n\in\llbracket 0,P\rrbracket$ assume the existence of a density $\mu_{n}(z)$ w.r.t. the Lebesgue measure. Assuming that $\psi_{n,k}$ is “volume preserving”, then for $k\in\llbracket 0,T\rrbracket$,
@@
When $\{\psi_{n,k},k\in\llbracket P\rrbracket\}$ are not volume preserving, additional multiplicative Jacobian determinant-like terms may be required (see Lemma 10 and the additional requirement that $\{\psi_{n,k},k\in\llbracket P\rrbracket\}$ be differentiable).
A central point throughout this paper is how samples from $\mu_{n}$ can be used to obtain samples from the marginal $\bar{\mu}_{n}$ and vice-versa, thanks to the mixture structure relating the two distributions. Naturally, given $z\sim\mu_{n}$, sampling $k\sim\mathcal{U}\big{(}\llbracket 0,T\rrbracket\big{)}$ and returning $\psi_{n,k}^{-1}(z)$ yields a sample from the marginal $\bar{\mu}_{n}$. More importantly, we have
Proposition 3 (Knitting-Tinking”).
For $n\in\llbracket 0,P\rrbracket$,
  1. 1.
    Assume $z\sim\bar{\mu}_{n}$ and sample $k\sim\bar{\mu}_{n}(\cdot\mid z)$. Then $z\mid k\sim\mu_{n}^{\psi_{n,k}^{-1}}({\rm d}z)$ and $\psi_{n,k}(z)\sim\mu_{n}$.
  2. 2.
    For $f\colon\mathsf{Z}\rightarrow\mathbb{R}$ $\mu_{n}$-integrable and $k\in\llbracket 0,T\rrbracket$
    $\int f(z)\mu_{n}({\rm d}z)$ =$=\int\Big{\{}\sum_{k=0}^{T}f\circ\psi_{n,k}(z)\bar{\mu}_{n}(k\mid z)\Big{\}}\bar{\mu}_{n}({\rm d}z)\,,$ (10)
Proof.
For $f\colon\mathsf{Z}\rightarrow\mathbb{R}$ $\mu_{n}$-integrable and $k\in\llbracket 0,T\rrbracket$, a change of variable (see Theorem 9) yields
@@
Therefore
@@
$=\sum_{k=0}^{T}\int f\circ\psi_{n,k}(z)\bar{\mu}_{n}(k,{\rm d}z)$
=$=\int\Big{\{}\sum_{k=0}^{T}f\circ\psi_{n,k}(z)\bar{\mu}_{n}(k\mid z)\Big{\}}\bar{\mu}_{n}({\rm d}z)\,,$
Using $f=\mathbf{1}_{A}$ for $A\in\mathscr{Z}$ formally establishes the earlier claim that if $(k,z)\sim\bar{\mu}_{n}$ then $\psi_{n,k}(z)\sim\mu_{n}$. ∎
Note that, as suggested by Example 2, construction of the estimator corresponding to (10) will only require evaluations of the density $\mu_{n}$ and function $f$ at $z,\psi_{n,1}(z),\psi_{n,2}(z),\ldots,\psi_{n,T}(z)$. It should be clear that the results of Proposition 3 extend well beyond the leapfrog integrator setup, a fact we discuss and exploit in the remainder of the manuscript.
We now turn to the description of an SMC algorithm targeting $\big{\{}\bar{\mu}_{n},n\in\llbracket 0,P\llbracket\big{\}}$, Alg. 3, and then establish that it is probabilistically equivalent to Alg. 2, in a sense made precise below. With $z=(x,v)$ for $n\in\llbracket P\rrbracket$ we introduce the following mutation kernel
@@
where we note that the refreshment kernel has the property that $\mu_{n-1}R_{n}=\mu_{n-1}$. With the assumption $\mu_{n-1}\gg\bar{\mu}_{n}$, from Lemma 13 one can identify the corresponding near optimal kernel $\bar{L}_{n-1}\colon\mathsf{Z}\times\mathscr{Z}\rightarrow[0,1]$, which leads to the SMC sampler importance weights, at step $n\in\llbracket P\rrbracket$,
@@
that is for $f\colon\mathsf{Z}\rightarrow\mathbb{R}$ such that $\bar{\mu}_{n}(|f|)<\infty$,
$\int f(z^{\prime})\frac{{\rm d}\bar{\mu}_{n}}{{\rm d}\mu_{n-1}}(z^{\prime})\bar{\mu}_{n-1}\otimes\bar{M}_{n}\big{(}{\rm d}(z,z^{\prime})\big{)}=\bar{\mu}_{n}(f)\,.$
@@
Notice that we assume here $\psi_{0}={\rm Id}$, hence that $\bar{\mu}_{0}=\mu_{0}$, and that the weights $w_{n,k}$ appear as being computed twice in step 3 and step 14 when evaluating the resampling weights at the previous iteration, for the only reason that it facilitates exposition. The identities (10) and (9) suggest, for any $n\in\llbracket P\rrbracket$, the estimator of $\mu_{n}(f)$ for $f\colon\mathsf{Z}\rightarrow\mathbb{R}$ $\mu_{n}$-integrable,
@@
a form of Rao-Blackwellization implying lower variance for $\hat{\mu}_{n}(f)$ while the two estimators share the same bias. Interestingly a result we establish later in the paper, Proposition 5, suggests that the variance $\check{\mu}_{n}(f)$ is smaller than that of the standard Monte Carlo estimator that assumes $N$ samples $z_{n}^{(i)}\overset{{\rm iid}}{\sim}\mu_{n}$, due to the control variate nature of integrator snippets estimators. Another point is that computation of the weight $\mu_{n}\circ\psi_{n,k}/\bar{\mu}_{n}$ only requires knowledge of $\mu_{n}$ up to a normalizing constant, that is the estimator is unbiased if $\check{z}_{n}^{(i)}\sim\bar{\mu}_{n}$ for $i\in\llbracket N\rrbracket$ even if $\mu_{n}$ is not completely known, while the estimator (7) will most often require self-normalisation, hence inducing a bias.
We now provide the probabilistic argument justifying Alg. 2 and the shared notation $\{z_{n}^{(i)},i\in\llbracket N\rrbracket\}$ in Alg. 2 and Alg. 3. The result is illustrated in Fig. 3.